OpenStudy (anonymous):

the length of a small rectangular room is 4 units more than the width, the area of the room is 32 square units. what is the dimensions of the room?

OpenStudy (anonymous):

Let, the length be x. So, the width is 4x. Now, area=length*width So, 32=4x*x 32=4x^2 x^2=8 x=2*rootover2 length=2*rootover2 width=8*rootover2

OpenStudy (anonymous):

Let the width=x and the length=x+4 Area=Length X Width The easy way to determine the two numbers you want is by making a list of the factors of 32. Here is mine. 1 X 32 2 X 16 4 X 8 Now when you look at those three equations, you can see which one matches your equation. In this case it is 4 X 8, since 8 is 4 more than 4. Your answer would be: Length=8 and Width=4